# Diffusive Gibbs Sampling

### v2

@@ Abstract
The inadequate mixing of conventional Markov Chain Monte Carlo (MCMC) methods for multi-modal distributions presents a significant challenge in practical applications such as Bayesian inference and molecular dynamics. Addressing this, we propose Diffusive Gibbs Sampling (DiGS), an innovative family of sampling methods designed for effective sampling from distributions characterized by distant and disconnected modes. DiGS integrates recent developments in diffusion models, leveraging Gaussian convolution to create an auxiliary noisy distribution that bridges isolated modes in the original space and applying Gibbs sampling to alternately draw samples from both spaces. Our approach exhibits a better mixing property for sampling multi-modal distributions than state-of-the-art methods such as parallel tempering. We demonstrate that our sampler attains substantially improved results across various tasks, including mixtures of Gaussians, Bayesian neural networks and molecular dynamics.
@@ 1 Introduction
$p(x)=\frac{\exp(-E(x)}{Z},$ (1)
where $x\in\mathbb{R}^{d}$ is the variable to be sampled, $E:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a lower-bounded differentiable energy function, and $Z=\int\exp(-E(x))\mathop{}\!\mathrm{d}{x}$ is the (intractable) normalization constant. Our goal is to draw independent samples $x\sim p(x)$ from the target distribution and estimate expectations of functions $\int h(x)p(x)\mathop{}\!\mathrm{d}{x}$ under the target distribution $p(x)$. Since $Z$ is independent of $x$ and $E$ is assumed to be differentiable, we can directly evaluate the gradient of the log density of the target distribution
@@
which is also known as the score function. This assumption is commonly satisfied in various practical applications, such as posteriors in Bayesian inference (Welling & Teh, 2011), score/energy networks in generative image modelling (Song & Ermon, 2019), and Boltzmann distributions in statistical mechanics (Noé et al., 2019). Various score-based sampling methods have been proposed to sample from unnormalized distributions, which we introduce below.
@@ 1.1 Score-Based MCMC Methods
Despite the introduction of HMC to address the challenge of exploring a distribution’s support, it can still be ineffective for distributions with disconnected modes in practice. This ineffectiveness is due to HMC’s dynamics struggling to cross low-density barriers that separate the modes, leading to prolonged transitions from one mode to another (Pompe et al., 2020). In general, score-based methods utilize local gradient information to propose subsequent states, which presents challenges in multi-modal distributions when there is insufficient bridging density to connect different modes. This limited connectivity impairs the resulting samples’ ability to accurately represent the entire distribution, a limitation illustrated in Figure 1. To enhance mode coverage in multi-modal sampling by bridging disconnected modes, a popular approach is Gaussian convolution, which has been widely used in the recent developments of diffusion models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020) which we discuss below.
@@ 1.2 Convolution-Based Method
For a target distribution $p(x)$ and a Gaussian convolution kernel $p(\tilde{x}|x)=\mathcal{N}(\tilde{x}|\alpha x,\sigma^{2}I)$, a convolved distribution $p(\tilde{x})$ can be constructed as follows:
@@
Since $p(\tilde{x}|x)$ has the full support of $\mathbb{R}^{d}$, it can effectively create non-negligible density paths between disconnected modes in $p(\tilde{x})$, which makes the modes in $p(\tilde{x})$ exhibit better connectivity compared to those in $p(x)$. This cherished property has made Gaussian convolution a popular remedy to heal the blindness of score matching (Song & Ermon, 2019; Wenliang & Kanagawa, 2020; Zhang et al., 2022) or fix KL divergence training for distributions with disjoint or ill-defined density (Roth et al., 2017; Zhang et al., 2020, 2023b; Brown et al., 2022).
Due to the mode-bridging property of convolution, it is generally easier for score-based samplers to explore the whole space of $p(\tilde{x})$ than $p(x)$. If we could obtain numerous samples $\tilde{x}\sim p(\tilde{x})$, then it is more likely that these samples will encapsulate a broader range of modes in $p(\tilde{x})$ which are close to different high-density areas in $p(x)$. Consequently, these samples of $\tilde{x}$ can then serve as initial points for sampling from the original target $p(x)$, which facilitates score-based samplers in capturing different modes in $p(x)$. However, the score function of the convolved distribution
@@
is typically intractable for non-Gaussian targets $p(x)$, which makes score-based sampling from $p(\tilde{x})$ infeasible.
In the next section, we introduce a novel sampling method named Diffusive Gibbs Sampling (DiGS), which leverages convolution to enhance multi-modal sampling while avoiding the intractability of the convolved score function.
@@ 2 Diffusive Gibbs Sampling
Instead of trying to directly produce samples from the intractable convolved distribution $p(\tilde{x})$, Diffusive Gibbs Sampling (DiGS) uses a Gibbs sampler to sample from the joint distribution $p(x,\tilde{x})=p(\tilde{x}|x)p(x)$. This Gibbs sampling procedure involves alternately drawing samples from the two conditional distributions $p(\tilde{x}|x)$ (the convolution kernel) and $p(x|\tilde{x})$ (the denoising posterior). In each step, for a given data sample $x^{(i-1)}$ from the target $p(x)$, we draw
@@
    a new data sample $x^{(i)}\sim p(x|\tilde{x}=\tilde{x}^{(i-1)})$.
For a Gaussian convolution kernel $p(\tilde{x}|x)=\mathcal{N}(\tilde{x}|\alpha x,\sigma^{2}I)$, noisy samples can be easily obtained by corrupting data samples with Gaussian noises:
$\tilde{x}=\alpha x+\sigma\epsilon,\quad\epsilon\sim N(0,I).$ (10)
Here, $10<\alpha<$ 1 is a contraction factor inspired by Diffusion models (Ho et al., 2020) and $\sigma$ determines the smoothness level of the Gaussian convolution. Intuitively, a small $\alpha$ will compress the distribution to make the modes closer and a large $\sigma$ will encourage the sampler to jump out of the local modes. In Section 2.2, we will examine the effects of these two parameters. Strategies for selecting these parameters in practical scenarios will be further discussed in Section 2.3. The denoising posterior $p(x|\tilde{x})$ is proportional to the joint distribution $p(x,\tilde{x})=p(\tilde{x}|x)p(x)$, taking the form
@@
which has a tractable score function (Gao et al., 2020; Huang et al., 2023):
@@
Therefore, common score-based methods like ULA, MALA, or HMC could be directly applied to sample from the denoising posterior $p(x|\tilde{x})$. It is worth noting that, in contrast to directly sampling from the original target $p(x)\propto\exp(-E(x))$ using score-based methods, incorporating an additional quadratic term as in Equation 11 improves the Log-Sobolev conditions, which in turn significantly increases the convergence speed of score-based samplers such as ULA (Vempala & Wibisono, 2019), see Huang et al. (2023) for further in-depth analysis. Below, we show that the DiGS yields a $p(x,\tilde{x})$-irreducible and recurrent Markov Chain under certain regularity conditions. Intuitively, these properties ensure that the chain comprehensively explores the state space from any starting point (irreducibility) and effectively captures the target distribution by infinitely revisiting every state (recurrence) (Robert et al., 1999).
@@
Figure 2: Visualization of an MoG target with unequal weights $w=[0.1,0.1,0.1,0.7]$ for different components. (a) Density heatmap of the target $p(x)$, a data sample $x^{(i-1)}$ and a noisy sample $\tilde{x}^{(i-1)}$. (b) Density heatmap of the denoising posterior $p(x|\tilde{x}^{(i-1)})$ with Gaussian convolution parameters $\alpha=1,\sigma=1$.
@@
The proof of Theorem 2.1 can be found in Appendix A. In addition to being a valid MCMC sampler, there are some practical considerations that can impact the performance of DiGS. We will discuss these in the following sections.
@@ 2.1 Initialization of the Denoising Sampling Step
Ideally, one might hope to construct an exact Gibbs sampler where the score-based sampler targeting $p(x|\tilde{x}^{(i-1)})$ draws a true sample $x^{(i)}$ at each iteration. Unfortunately, when the target distribution has very disconnected modes, the resulting posterior $p(x|\tilde{x})$ may still exhibit a multi-modal nature. For instance, Figure 2(a) illustrates the density of an MoG with unbalanced weights. For a given previous data sample $x^{(i-1)}$, we generate a noisy sample $\tilde{x}^{(i-1)}$ using a Gaussian convolutional kernel with $\alpha=1$ and $\sigma=1$. The corresponding denoising density $p(x|\tilde{x}^{(i-1)})$ is depicted in Figure 2(b), which exhibits four distinct modes with varying weights. In such scenarios, selecting an appropriate initial point, $x^{(i)}_{init}$, for the subsequent sampling process $x^{(i)}\sim p(x|\tilde{x}=\tilde{x}^{(i-1)})$ is crucial for score-based samplers.
@@
Figure 3: Comparison of different initialization techniques for denoising posterior sampling on an unequal weighted MoG target described in Figure 2(a). In each case, we generate 1,000 samples using a Gaussian convolution kernel with $\alpha=1,\sigma=1$.
Table 1: MMD between true samples and samples obtained using different initialization techniques for denoising posterior sampling on an unequal weighted MoG target described in Figure 2(a).
@@
Figure 4: Effects of the hyperparameters $\alpha,\sigma$ in Gaussian convolution kernels and the number $T$ of noise levels in the variance-preserving (VP) noise scheduling. The y-axis in all three plots is the MMD between true samples and generated samples generated by DiGS with varying hyperparameters. The x-axis is $\alpha$ in (a), $\sigma$ in (b), and $T$ in (c). Experimental setups can be found in Appendices C.2 and C.3.
An ideal initial point for sampling from the denoising posterior $p(x|\tilde{x}^{(i-1}))$ would be the mean of the denoising distribution, defined as $\mu(\tilde{x})\equiv\int xp(x|\tilde{x})\mathop{}\!\mathrm{d}{x}$. By Tweedie’s lemma (Efron, 2011; Robbins, 1992), the mean can be expressed as a function of the score $\nabla_{\tilde{x}}\log p(\tilde{x})$:
Alternatively, a straightforward approach is to initialize the denoising sampler at the previous state $x^{(i-1)}$. However, this strategy has a drawback: since $x^{(i-1)}$ is typically close to one of the modes, the score-based sampler often remains trapped in the vicinity of that mode, thereby hindering effective exploration of the entire distribution. Another heuristic initialization strategy is to use $\tilde{x}^{(i-1)}/\alpha$, where the scaling factor $\alpha$ reflects the scale relationship between $x$ and $\tilde{x}$, as suggested by the mean function in Equation 13. This approach exhibits a uniform preference for a random mode (for instance, the upper-left mode in Figure 2(b)) and ignores the underlying weighting between modes, resulting in a bias in representing the true weights of different modes and consequently diminishing the overall quality of the samples.
To avoid potential bias, rather than aiming to draw an exact sample from $p(x|\tilde{x}^{(i-1)})$ at each iteration we employ a Metropolis-within-Gibbs scheme, where a finite number of score-based MCMC updates are applied in each sweep. We propose to facilitate mixing across modes by sampling an initialization position $x_{init}^{\prime}$ for the subsequent score-based sampling steps from an additional MCMC transition kernel. Specifically, to initialize denoising sampling for the posterior $p(x|\tilde{x}^{(i-1)})$, we employ a proposal $x_{init}^{\prime}\sim q(x|\tilde{x}^{(i-1)})$ which is centered at the (scaled) noisy sample $\tilde{x}^{(i-1)}/\alpha$:
@@
where the mean and the variance are inspired by the mean function $\mu(\tilde{x}^{(i-1)})$. Note that this proposal only depends on the noisy sample $\tilde{x}^{(i-1)}$ and is independent of the previous state $x^{(i-1)}$. We use the MH algorithm to calculate the acceptance rate for this proposal:
@@
To demonstrate the benefits of the additional MCMC kernel updating the initialization, we run DiGS with three different initializations – the previous state $x^{(i-1)}$, a heuristic re-initialization at $\tilde{x}^{(i-1)}/\alpha$, and our MH transition strategy – on an MoG target with different component weights. Figure 3 provides a visual comparison of the samples obtained from these initializations, showing that only the MH transition scheme captures all modes with the correct weightings. For evaluation, we employ the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) throughout all the experiments. The results are shown in Table 1, which demonstrates that the MH strategy outperforms the other two in capturing all modes and accurately representing the true weightings. Detailed experimental setup can be found in Appendix C.1.
@@
1:  Input: Target energy $E(x)$; Gaussian convolution hyperparameters $\alpha,\sigma$; score-based denoising sampler $\mathcal{S}$; the number of denoising sampling steps $L$; the number of Gibbs sampling sweeps $K$; initial data sample $x^{(0)}$.
@@
6:     Draw $x^{(i)}\sim p(x|\tilde{x}^{(i-1)})$ by running the score-based sampler $\mathcal{S}$ for $L$ steps from the initial point $x^{(i)}_{init}$.

### v5

@@ Abstract
The inadequate mixing of conventional Markov Chain Monte Carlo (MCMC) methods for multi-modal distributions presents a significant challenge in practical applications such as Bayesian inference and molecular dynamics. Addressing this, we propose Diffusive Gibbs Sampling (DiGS), an innovative family of sampling methods designed for effective sampling from distributions characterized by distant and disconnected modes. DiGS integrates recent developments in diffusion models, leveraging Gaussian convolution to create an auxiliary noisy distribution that bridges isolated modes in the original space and applying Gibbs sampling to alternately draw samples from both spaces. A novel Metropolis-within-Gibbs scheme is proposed to enhance mixing in the denoising sampling step. DiGS exhibits a better mixing property for sampling multi-modal distributions than state-of-the-art methods such as parallel tempering, attaining substantially improved performance across various tasks, including mixtures of Gaussians, Bayesian neural networks and molecular dynamics.
@@ 1 Introduction
$p(x)=\frac{\exp(-E(x))}{Z},$ (1)
where $x\in\mathbb{R}^{d}$ is the variable of interest, $E:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a lower-bounded differentiable energy function, and $Z=\int\exp(-E(x))\mathop{}\!\mathrm{d}{x}$ is the (intractable) normalization constant. We aim to draw independent samples $x\sim p(x)$ from the target distribution and estimate expectations of functions $\mathbb{E}_{p(x)}[h(x)]=\int h(x)p(x)\mathop{}\!\mathrm{d}{x}$ under the target distribution $p(x)$.
The gradient of the log density of the target distribution is known as the score function:
@@
which is independent of $Z$. The score function can be evaluated at any location $x$ since $E$ is assumed to be differentiable. This assumption is commonly satisfied in various practical applications, such as posteriors in Bayesian inference (Welling & Teh, 2011), score/energy networks in generative image modelling (Song & Ermon, 2019), and Boltzmann distributions in statistical mechanics (Noé et al., 2019). Below, we introduce score-based methods for sampling from unnormalized distributions.
@@ 1.1 Score-Based MCMC Methods
Despite the effort to address the challenge of exploring the whole support of a distribution using Langevin and Hamiltonian dynamics, MALA and HMC can still be ineffective for distributions with disconnected modes in practice since they often struggle to cross low-density barriers that separate the modes, leading to prolonged transitions from one mode to another (Pompe et al., 2020). More generally, score-based sampling methods utilize local gradient information to propose subsequent states, which presents challenges in sampling multi-modal distributions when there is insufficient bridging density to connect different modes. This limited connectivity impairs the ability of the resulting samples to accurately represent the entire distribution, a limitation illustrated in Figure 1.
@@ 1.2 Convolution-Based Method
A popular approach to bridging disconnected modes in multi-modal distributions is Gaussian convolution, which has been widely used in the recent developments of diffusion models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020). For a target distribution $p(x)$ and a Gaussian convolution kernel $p(\tilde{x}|x)=\mathcal{N}(\tilde{x}|\alpha x,\sigma^{2}I)$, a convolved distribution $p(\tilde{x})$ can be constructed as follows:
@@
Since $p(\tilde{x}|x)$ has the full support of $\mathbb{R}^{d}$, it can effectively create non-negligible density paths between disconnected modes in $p(\tilde{x})$, which makes the modes in $p(\tilde{x})$ exhibit better connectivity compared to those in $p(x)$. This cherished property has made Gaussian convolution a popular remedy to heal the blindness of score matching (Song & Ermon, 2019; Wenliang & Kanagawa, 2020; Zhang et al., 2022) and fix KL divergence training for distributions with disjoint or ill-defined density (Roth et al., 2017; Zhang et al., 2020, 2023b; Brown et al., 2022).
Due to the mode-bridging property of convolution, it is generally easier for score-based samplers to explore the whole space of $p(\tilde{x})$ than $p(x)$. If we could obtain numerous samples $\tilde{x}\sim p(\tilde{x})$, then it is more likely that these samples will encapsulate a broader range of modes in $p(\tilde{x})$ which are close to different high-density areas in $p(x)$. Consequently, these samples of $\tilde{x}$ can then serve as initial points for sampling from the original target $p(x)$, which facilitates score-based samplers in capturing different modes in $p(x)$. However, the score function of the convolved noisy distribution $p(\tilde{x})$ has the form
@@
which is typically intractable for non-Gaussian targets. This makes score-based sampling infeasible for $p(\tilde{x})$.
@@ 2 Diffusive Gibbs Sampling
We introduce a novel sampling method, Diffusive Gibbs Sampling (DiGS), which leverages Gaussian convolution with an innovative Metropolis-within-Gibbs scheme to enhance multi-modal sampling, while avoiding the intractability of the convolved score function as shown in Equation 9.
2.1 Sampler Construction
Instead of trying to directly produce samples from the intractable convolved distribution $p(\tilde{x})$, DiGS employs a Gibbs sampler to sample from the joint distribution $p(x,\tilde{x})=p(\tilde{x}|x)p(x)$. This Gibbs sampling procedure involves alternately drawing samples from the two conditional distributions $p(\tilde{x}|x)$ (the convolution kernel) and $p(x|\tilde{x})$ (the denoising posterior). In each step, for a given clean sample $x^{(i-1)}$ from the target $p(x)$, we draw
@@ 2.1 Sampler Construction
    a new clean sample $x^{(i)}\sim p(x|\tilde{x}=\tilde{x}^{(i-1)})$.
For a Gaussian convolution kernel $p(\tilde{x}|x)=\mathcal{N}(\tilde{x}|\alpha x,\sigma^{2}I)$, noisy samples can be easily obtained by corrupting clean samples with Gaussian noises:
$\tilde{x}=\alpha x+\sigma\epsilon,\quad\epsilon\sim N(0,I),$ (10)
where $\alpha\in[0,1]$ is a contraction factor inspired by Diffusion models (Ho et al., 2020) and $\sigma$ determines the level of smoothness in the Gaussian convolution. Intuitively, a small $\alpha$ will compress the distribution and make the modes closer, and a large $\sigma$ will encourage the sampler to jump out of the local modes. The effects of these hyperparameters will be investigated in Section 2.3, and hyperparameter tuning strategies will be further discussed in Section 2.4.
Unlike the convolved distribution $p(\tilde{x})$ which has an intractable score as shown in Equation 9, the denoising posterior $p(x|\tilde{x})\propto p(x,\tilde{x})=p(\tilde{x}|x)p(x)$, taking the form
@@
has a tractable score function (Gao et al., 2020; Huang et al., 2023):
@@
Therefore, common score-based methods like MALA, and HMC could be directly applied to sample from the denoising posterior $p(x|\tilde{x})$ in the denoising sampling step. It is worth noting that, in contrast to directly sampling from the original target $p(x)\propto\exp(-E(x))$ using score-based methods, incorporating an additional quadratic term as in Equation 11 improves the Log-Sobolev conditions (i.e., makes it more “Gaussian-like”), which in turn significantly increases the convergence speed of score-based samplers (Vempala & Wibisono, 2019); see Huang et al. (2023) for further in-depth analysis. Below, we show that the DiGS yields a $p(x,\tilde{x})$-irreducible and recurrent Markov Chain under certain regularity conditions.
@@
Figure 2: Visualization of an MoG target with unequal weights $w=[0.1,0.1,0.1,0.7]$ for different components. (a) Density heatmap of the target $p(x)$, a clean sample $x^{(i-1)}$ and a noisy sample $\tilde{x}^{(i-1)}$. (b) Density heatmap of the denoising posterior $p(x|\tilde{x}^{(i-1)})$ with Gaussian convolution parameters $\alpha=1,\sigma=1$.
@@
The proof of Theorem 2.1 can be found in Appendix A. Intuitively, these properties ensure that the chain comprehensively explores the state space from any starting point (irreducibility) and effectively captures the target distribution by infinitely revisiting every state (recurrence) (Robert et al., 1999).
In addition to being a valid MCMC sampler, there are some practical considerations that can impact the performance of DiGS. We discuss these in the following sections.
@@ 2.2 Initialization of the Denoising Sampling Step
Ideally, one might hope to construct an exact Gibbs sampler where the score-based sampler targeting $p(x|\tilde{x}^{(i-1)})$ draws a true sample $x^{(i)}$ at each iteration. Unfortunately, when the target distribution has very disconnected modes, the resulting denoising posterior $p(x|\tilde{x})$ may still exhibit a multi-modal nature. For instance, Figure 2(a) illustrates the density of an MoG with unbalanced weights. For a given previous clean sample $x^{(i-1)}$, we generate a noisy sample $\tilde{x}^{(i-1)}$ using a Gaussian convolutional kernel with $\alpha=1$ and $\sigma=1$. The corresponding denoising density $p(x|\tilde{x}^{(i-1)})$ is depicted in Figure 2(b), which exhibits four distinct modes with varying weights. In such scenarios, selecting an appropriate initial point, $x^{(i)}_{init}$, for the subsequent sampling process $x^{(i)}\sim p(x|\tilde{x}=\tilde{x}^{(i-1)})$ is crucial for score-based samplers.
An ideal initial point for sampling from the denoising posterior $p(x|\tilde{x}^{(i-1})$ would be the mean of the denoising distribution, defined as $\mu(\tilde{x})\equiv\int xp(x|\tilde{x})\mathop{}\!\mathrm{d}{x}$. By Tweedie’s lemma (Efron, 2011; Robbins, 1992), the mean can be expressed as a function of the noisy score $\nabla_{\tilde{x}}\log p(\tilde{x})$:
Alternatively, a more straightforward approach is to initialize the denoising sampler at the previous state $x^{(i-1)}$. However, this strategy has a drawback: since $x^{(i-1)}$ is typically close to one of the modes, the score-based sampler often remains trapped in the vicinity of that mode, thereby hindering effective exploration of the entire distribution. Another heuristic initialization strategy is to use the (scaled) noisy sample $\tilde{x}^{(i-1)}/\alpha$, where the scaling factor $\alpha$ reflects the scale relationship between $x$ and $\tilde{x}$ as suggested by the mean function in Equation 13. This approach exhibits a uniform preference for a random mode (for instance, the upper-left mode in Figure 2(b)) and ignores the underlying weighting between modes, resulting in a bias in representing the true weights of different modes and consequently diminishing the overall quality of the samples.
@@
Figure 3: Comparison of different initialization techniques for denoising posterior sampling on an unequally weighted MoG target described in Figure 2(a). In each case, we generate 1,000 samples using a Gaussian convolution kernel with $\alpha=1,\sigma=1$.
Table 1: MMD between true samples and samples obtained using different initialization techniques for denoising posterior sampling on an unequally weighted MoG target described in Figure 2(a).
@@
Figure 4: Effects of the hyperparameters $\alpha,\sigma$ in Gaussian convolution kernels and the number $T$ of noise levels in the variance-preserving (VP) noise scheduling. The y-axis in all three plots is the MMD between true samples and generated samples generated by DiGS with varying hyperparameters. Experimental setups can be found in Appendices D.2 and D.3.
2.2.1 A Metropolis-within-Gibbs Scheme
To avoid potential bias, we propose a Metropolis-within-Gibbs scheme to facilitate mixing across modes by sampling an initialization position $x_{init}^{\prime}$ from an additional MCMC transition kernel for the subsequent finite score-based sampling steps. Specifically, to initialize score-based sampling from the denoising posterior $p(x|\tilde{x}^{(i-1)})$, we employ a linear Gaussian proposal $x_{init}^{\prime}\sim q(x|\tilde{x}^{(i-1)})$ which is centered at the (scaled) noisy sample $\tilde{x}^{(i-1)}/\alpha$:
@@ 2.2.1 A Metropolis-within-Gibbs Scheme
where the mean and the variance are chosen according to $q(x|\tilde{x})\propto p(\tilde{x}|x)$ inspired by the mean function $\mu(\tilde{x}^{(i-1)})$. Note that this proposal only depends on the noisy sample $\tilde{x}^{(i-1)}$ and is independent of the previous state $x^{(i-1)}$. We use the MH algorithm to calculate the acceptance rate for this proposal:
@@
To demonstrate the benefits of the additional MCMC kernel updating the initialization, we run DiGS with three different initializations (the previous state $x^{(i-1)}$, a heuristic re-initialization at $\tilde{x}^{(i-1)}/\alpha$, and our MH transition strategy) on an MoG target with different component weights. Figure 3 provides a visual comparison of the samples obtained from these initializations, showing that only the MH transition scheme captures all modes with the correct weightings. For evaluation, we employ the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) throughout all the experiments. The results are shown in Table 1, which demonstrates that the MH strategy outperforms the other two in capturing all modes and accurately representing the true weightings. Detailed experimental setup can be found in Appendix D.1.
@@
1:  Input: target energy $E(x)$; Gaussian convolution hyperparameters $\alpha,\sigma$; score-based denoising sampler $\mathcal{S}$; the number of denoising sampling steps $L$; the number of Gibbs sampling sweeps $K$; initial clean sample $x^{(0)}$.
@@
6:     Draw $x^{(i)}\sim p(x|\tilde{x}^{(i-1)})$ by running the score-based sampler $\mathcal{S}$ for $L$ steps from the initial point $x^{(i)}_{init}$ using Equation 12.